# Optimal sensing of photon addition and subtraction on nonclassical light

### v1

@@ IV Photon Subtracted Squeezed Vacuum States
Photon subtraction from the SVS is carried out in theory by repeated application of $\hat{a}$ on the SVS and suitably normalizing the resulting state. The $1$-photon subtracted SVS
$|\xi,-1\rangle=\mathcal{K}_{-1}\sum_{n=0}^{\infty}(-1)^{n}\frac{\sqrt{(2n+1)}!}{2^{n}n!}e^{in\phi}\tanh^{n}r~{}|2n+1\rangle,$ (11)
with $\mathcal{K}_{-1}=-e^{i\phi}(\cosh r)^{-3/2}$. The $2$-photon subtracted SVS
$|\xi,-2\rangle=\mathcal{K}_{-2}\sum_{n=2}^{\infty}\frac{(-1)^{n}}{\sqrt{(2n-2)!}}\frac{(2n)!}{2^{n}n!}e^{in\phi}\tanh^{n}r~{}|2n-2\rangle,$ (12)
[…]
In Fig. 4 we consider the plot of $W_{1}$ between |$\ket{\xi}$ and |$\ket{\xi,-1}$ (equivalently |$\ket{\xi,1}$) versus $r$ (black circles). When compared with $W_{1}$ between |$\ket{\xi}$ and |$\ket{\xi,-2}$ (green triangles), we see that the two distances are equal at $r\approx 0.66$, beyond which there is a crossover, and the trends in $W_{1}$ shift. As in the case of photon addition, such crossovers are not seen in other quadratures (Fig. 4 in the SM).
@@ V Concluding remarks
In continuous variable systems, quantum states could suffer from errors arising through photon loss or gain. A precise estimate of the number of such photons is a prerequisite for error correction. High precision detection is possible with states with large sensitivity or small variance in those quadratures where the changes arise due to such errors [50]. We have shown that, for appropriate ranges of the state parameter, the effect of one or two photon addition can be captured in the Wasserstein distances computed between the nonclassical reference state (the squeezed vacuum, for instance) and its photon added counterparts. Similar results hold in the case of the even cat state over an appropriate range of parameter values. It is worth noting that noise estimation gets enhanced by addition of photons to the original nonclassical state [5, 51].
Our approach is based on examining the qualitative features of the relevant optical tomograms (histograms of experimental data). In particular we have examined the shifts in the intensities of nonclassical light in specific quadratures, arising due to photon addition or subtraction. The PDFs comprising the relevant tomograms corroborate our findings. We have demonstrated the usefulness of this tomographic approach both for identifying qualitative changes and for quantifying such changes using the Wasserstein distance. In particular, we have identified specific small ranges of values of state parameters (the crossovers) which are not sufficiently sensitive to photon addition to, or removal from the nonclassical reference state. Furthermore, we have also obtained a sufficiently wide range of experimentally accessible parameter values where it is possible to quantify with precision the extent of change in the photon number by computing the Wasserstein distance directly from the tomograms.
In conclusion, we have presented an experimentally optimal, tomogram-based measure of quantum state discrimination for sensing subtle differences between nonclassical states. It is based on the Wasserstein distance, a true metrological measure, that discriminates states with addition and subtraction of photons. Its advantages are illustrated through examples of nonclassical states, ranging from squeezed vacuum to cat states and their generalizations. Our approach suggests the possibility of sensing changes in the photon number directly from the optical tomogram, circumventing detailed state reconstruction. This could be potentially useful for identifying such changes in generic multimode states where reconstructing the Wigner function from the tomogram could pose considerable challenges.
@@ Acknowledgements.
We acknowledge partial support through funds from Mphasis to the Centre for Quantum Information, Computing and Communication (CQuICC), Indian Institute of Technology Madras. SL and VB thank the Department of Physics, Indian Institute of Technology Madras for infrastructural support.

### v2

@@ IV Photon Subtracted Squeezed Vacuum States
Photon subtraction from the SVS is carried out in theory by repeated application of $\hat{a}$ on the SVS and normalizing the resulting state. The $m$-photon subtracted SVS is given by
$|\xi,-m\rangle=\frac{1}{{\mathcal{K}^{1/2}_{-m}}}\sum_{n=\left\lceil m/2\right\rceil}^{\infty}\frac{(2n)!e^{in\phi}(-\tanh r)^{n}}{2^{n}n!\sqrt{(2n-m)!}}~{}|2n-m\rangle,$ (10)
where
$\mathcal{K}_{-m}=\sum_{n=\left\lceil m/2\right\rceil}^{\infty}\left[\frac{(2n)!}{2^{n}n!}\right]^{2}\frac{(\tanh r)^{2n}}{(2n-m)!}$ (11)
and $\left\lceil m/2\right\rceil$ denotes the smallest integer $\geqslant m/2$. It is clear from Eqs. (9) and (10) that one photon addition or subtraction operations on $|\xi\rangle$, are equivalent [24]. Tomograms corresponding to photon subtraction from the SVS are shown in Figs. 1 (b), (e) and (f) above. Their zoomed-in versions are given in SM (Figs. 1 (b), (e) and (f)).
Analogous to Fig. 3, we have also examined the crossovers in the plots of the Wasserstein distances versus $r$ between $|\xi\rangle$ and the different photon subtracted states $|\xi,-m\rangle$ ($m=1,2,3$). In the cases we have considered, there is no crossover along the $x$-quadrature. There are however, other quadratures, where such crossovers occur, as seen in Fig. 7 of SM. In both photon addition and subtraction, it is evident that the value of $r$ at a crossover depends on both the quadrature and the states considered. Hence, the neighborhood of crossovers should be avoided in order to see a clear distinction between states using the Wasserstein distance.
@@ V Concluding remarks
In continuous variable systems, quantum states could suffer from errors arising through photon loss or gain. A precise estimate of the number of such photons is a prerequisite for error correction. High precision detection is possible using states with large sensitivity or small variance in those quadratures where the changes arise due to such errors [70]. We have shown that, for appropriate ranges of the state parameter, the effect of photon addition or subtraction can be captured in the Wasserstein distances computed between the nonclassical reference state (the squeezed vacuum, for instance) and the corresponding photon added/subtracted states. We have further shown that similar results hold in the case of photon addition to the even cat state over an appropriate range of parameter values. It is worth noting that noise estimation gets enhanced by addition of photons to the original nonclassical state [5, 71].
Our approach is based on examining the qualitative features of the relevant optical tomograms (histograms of experimental data). In particular we have examined the shifts in the intensities of nonclassical light in specific quadratures, arising due to photon addition or subtraction. The PDFs comprising the relevant tomograms corroborate our findings. We have demonstrated the usefulness of this tomographic approach both for identifying qualitative changes and for quantifying such changes using the Wasserstein distance. In particular, we have identified specific small ranges of values of state parameters (crossover values) which are not sufficiently sensitive to photon addition to, or removal from the nonclassical reference state. We have also identified a sufficiently wide range of experimentally accessible parameter values where it is possible to quantify, with precision, the extent of change in the photon number by computing the Wasserstein distance directly from the tomograms.
In conclusion, we have presented an experimentally viable, tomogram-based measure of quantum state discrimination (the Wasserstein distance) for sensing subtle differences between nonclassical states. The Wasserstein distance is a true metrological measure, that discriminates between states that arise from the addition and subtraction of photons. Its advantages are illustrated through examples of nonclassical states, ranging from the squeezed vacuum to cat states and their generalizations. Our approach suggests the possibility of sensing changes in the photon number directly from the optical tomogram, circumventing detailed state reconstruction. This could be potentially useful for identifying such changes in generic multimode states where reconstructing the Wigner function from the tomogram could pose considerable challenges.
@@ Acknowledgements.
We acknowledge partial support through funds from Mphasis to the Centre for Quantum Information, Communication and Computing (CQuICC), Indian Institute of Technology Madras. SL and VB thank the Department of Physics, Indian Institute of Technology Madras for infrastructural support. We thank the referees for their valuable suggestions.